OpenStudy (anonymous):

-2 times the quantity q minus 3

OpenStudy (ddcamp):

Use the distributive property: \[a(b+c) = ab+ac\]

OpenStudy (the_fizicx99):

-2q + 6 :) -2*q = -2q -2 * (-3) = 6

OpenStudy (skullpatrol):